Find the following probabilities for a standard normal variable, Z
1) P(Z<-1.27) 2) P(-2.03<Z<3.49) 3) P(Z>1.74) 4)P(Z<0.17)

B. Find z if we know that the area to the left of z (under the normal curve) is 0.9265.
Answer #1

given the Z variable. is standard normal 2) P(Z<-1-27)=0.1020 #by using Z - table 10.07 -1.20 27 P(-2.03CZ < 3:49) = P(Z (3:447 P(Z <0.17)= 0.5675 #by using S z table. By The z value such that the left of z is 0.9265 obtained by! P(Z cz)= 0.9265 => z